1. a company has a current ratio of 3:1. which action will decrease this ratio?

a. issue of long term debentures
b. sale of equipment for cash
c. declaration of a dividend
d. collection of an accounts receivable

2. which price earning ratios and earnings yield do not match?

a. price earning ratio of 5; earnings yield of 20%
b. price earning ratio of 8; earnings yield of 12.5%
c. price earning ratio of 9; earnings yield of 12%
d. price earning ratio of 10; earnings yield of 10%
Answer #1

Solution: 1 :

The declaration of dividend will decrease the current ratio. The declaration of dividend will result in a Debit to the Retained earnings account and a corresponding credit to the Dividend payable account, with the amount of dividend declared.

Thus there will be an increase in the balance of the Dividend Payable account. The Dividend payable account as we know is a Current Liability. Hence increase in Dividend payable account will result in increase in current liability. This increase in current liability will decrease the current ratio.

Thus the solution is Option C. Declaration of a dividend.

The other options are incorrect due to the reasons given below as follows:

a.issue of long term debentures

The issue of long term debentures affects two accounts. It increase cash account by the amount of the long term debentures issued. Also, there is an increase in the Long term debt amount by the amount of long term debentures issued.

Thus there in Increase in Current asset i.e., Cash and a corresponding increase in long term liability. This increase in current assets will increase the current ratio.


b. sale of equipment for cash

The sale of equipment for cash affects two accounts. It increase cash account by the amount received on sale of equipment. Also, it decreases the Fixed asset balance.

Thus there in Increase in Current asset i.e., Cash and a corresponding decrease in Fixed assets. This increase in current assets will increase the current ratio.



d. collection of an accounts receivable

The collection of accounts receivable affects two accounts. It increase cash account by the amount of receivables collected. Also, it decreases the Accounts receivable balance.

Thus there in Increase in Current asset i.e., Cash and a corresponding decrease in Current assets i.e., accounts receivable of the same amount. Thus this transaction has no effect on the existing current ratio. The current ratio remains unchanged.

Solution : 2 :

The formula for calculating the Earnings yield based on price earnings ratio is

Earnings yield = 1 / ( P/E Ratio )

a. price earning ratio of 5; earnings yield of 20%

Based on the Price earning ratio of 5, Earnings yield = 1 / 5 = 0.20 = 20 %


b. price earning ratio of 8; earnings yield of 12.5%

Based on the Price earning ratio of 8, Earnings yield = 1 / 8 = 0.1250 = 12.50 %


c. price earning ratio of 9; earnings yield of 12%

Based on the Price earning ratio of 9, Earnings yield = 1 / 9 = 0.1111 = 11.11 %


d. price earning ratio of 10; earnings yield of 10%

Based on the Price earning ratio of 10, Earnings yield = 1 / 10 = 0.10 = 10 %

From the above calculations, it is clear that the Statement c. does not match. This is because a Price Earning ratio of 9 results in a Earnings Yield of 11.11 %